# Quenched coalescent for diploid population models with selfing and overlapping generations

### v1

@@ 7.3 Random walks on $\Theta$-valued processes
Note that $\Pi^{N,n}_{G}$ in Definition 7.9 is equal to $\Pi^{n}(\bar{G}^{N})$, where $\bar{G}^{N}$ is the sped-up ancestral graph of Lemma 7.8.
@@
Let $\mathcal{E}:=\bigsqcup_{n\in\mathbb{N}}\mathcal{E}_{n}$. The measure-valued map $\Phi:\mathcal{D}\left(\mathbb{R}_{+},\Theta\right)\to\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}\right)\right)$ defined by
@@
Suppose that $g_{N}$ is a sequence of elements in $\mathcal{D}\left(\mathbb{R}_{+},\Theta\right)$ that converge therein to an element $g$. We need to show that $\Phi(g_{N})$ converges weakly in $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}\right)\right)$ to $\Phi(g)$. That is, we need to show that for any dense subset $\mathcal{F}$ of functions $f$ in $C_{b}(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}\right))$ that $\int fd\Phi(g_{N})$ converges to $\int fd\Phi(g)$. As $\Theta$ is a locally compact Polish space, by Theorem 6.5 it suffices to check that
@@
for any collection $\left(\varphi_{i}\right)_{1\leq i\leq k}$ of elements in $C_{c}\left(\mathbb{R}_{+}\times\Theta\right)$.
As $g_{N}$ converges to $g$ and $\Theta$ is a discrete space we have that, for large enough $N$, that $g_{N}(0)$ is eventually constant. In particular, we may restrict to the case that $g_{N}(0)=(n,m_{0},\xi_{0})$ for all $N$. This allows us to consider continuity of the restriction of $\Phi$ as a map into $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$ without loss of generality.
As each $\varphi_{i}$ is compactly supported on $\mathbb{R}_{+}\times\Theta$, there is a $T>0$ such that $\cup_{i}supp(\varphi_{i})\subset[0,T]\times E$. Notice that, for any $N$ that
@@
The quantity on the right of (7.19) depends only on discrete graph structure of $g_{N}$, i.e. the finite sequence of values that $g_{N}$ takes, and the edge lengths of the graph $g_{N}$, i.e. the jump times. As $\Theta$ is discrete and $g_{N}\to g$ in the Skorokhod $J_{1}$ topology, for large enough $N$ the discrete graph structure of $g_{N}$ is eventually constant. Further, the edge lengths of $g_{N}$ converge uniformly on $[0,T]$ to those of $g$. This yields the desired convergence.
@@
As a corollary, we have that if $G_{N}$ is a sequence of $\mathcal{D}\left(\mathbb{R}_{+},\Theta\right)$-valued random variables converging weakly to $G$, then $\Phi(G_{N})$ converges weakly in distribution to $\Phi(G)$.
@@
To prove Theorem 4.7, we establish a lemma demonstrating the equivalence of the quenched coalescent for the $Q$-$\lambda$ graph, as in Definition 7.3, and the quenched coalescent for an EFC, as in Definition 4.6. Before the formal statement of the lemma, we give an intuitive explanation for why the two constructions agree. Fix any finite sample of at most $n$ particles and follow their coalescing random-walk trajectories. Under both the $Q$-$\lambda$ graph and the EFC, coalescences of the walks occur with the same structures and rates and fragmentations of each walk occur with the same rate. Because these rates uniquely determine the law of the induced coalescent on partitions and because both the $Q$-$\lambda$ graph and EFC give the same amount of information on the particle trajectories, the quenched coalescent laws coincide. The lemma that follows makes this argument rigorous via a joint coupling of the $Q$-$\lambda$ graph, the EFC, and random walks thereon.
@@
Suppose that $Q=\left(Q_{n}\right)_{n\in\mathbb{N}}$ is a consistent family of generators for a $\Xi$-coalescent for some $\Xi$ in $\mathcal{M}_{1}\left(\Delta\right)$. Let $\Xi^{\prime}=\Xi-\delta_{\mathbf{0}}\Xi\left(\mathbf{0}\right)$. Let $\Pi$ be an EFC process with characteristics $c_{k}=\Xi(\mathbf{0})$, $\nu_{Coag}=\nu_{\Xi^{\prime}}$, $c_{e}=\lambda$, and $\nu_{Disl}=0$. Then the random measures $\mathcal{L}_{\Pi}^{n}$ and $\mathcal{L}_{Q,\lambda}^{n}$ are equal in distribution.
@@
For a given $Q$-$\lambda$ graph $G=\left((l(t),m(t),\xi(t))\right)_{t\geq 0}$ and an EFC process $\Pi=\left(\Pi(t)\right)_{t\geq 0}$, we let $x^{G}=\left(x_{i}^{G}\right)_{1\leq i\leq n}$ and $x^{\Pi}=\left(x_{i}^{\Pi}\right)_{1\leq i\leq n}$ denote the family of coalescing random walks on $G$ and $\Pi$, respectively. Let $\chi^{n}_{G}$ and $\chi^{n}_{\Pi}$ denote the associated coalescents for $x^{G}$ and $x^{\Pi}$, respectively. We describe a coupling of $\Pi$, $G$, $x^{\Pi}$, and $x^{G}$ together so that the claim holds. Specifically, it will be such that $\chi^{n}_{G}=\chi^{n}_{\Pi}$ pointwise, and such that $\mathbb{P}\left(\chi^{n}_{\Pi}\in\cdot\,\mid\,\Pi\right)=\mathbb{P}\left(\chi^{n}_{\Pi}\in\cdot\,\mid\,G\right)$, which together prove the claim. The construction follows that for the EFC (Ber, 04, Section 3.2).
@@
      and if $k(t)\in\{x_{i}^{\Pi}(t)\}_{1\leq i\leq n}$, then $G(t)=\left(l(t-),m(t),\xi(t-)\right)$, where $m(t)$ is equal to $x_{i}^{G}(t-)$ (for any $i$ such that $x_{i}^{\Pi}(t-)=k(t)$; check that this is indeed well-defined); for $H$ an independent Bernoulli 010-1 random random variable with $\mathbb{P}\left(H=0\right)=\mathbb{P}\left(H=1\right)=\frac{1}{2}$ and if the only non-singleton block of $\xi^{(F)}(t)$ is $\{r,s\}$, then
@@
Suppose that, as $N\to\infty$, that $c_{N}\to 0$, $d_{N}c_{N}^{-1}\to\lambda<\infty$, and $\mathcal{L}^{N,n}_{G}$ converges in distribution in $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$ to $\mathcal{L}$. Then $\mathcal{L}^{N,n}$ also converges in distribution to $\mathcal{L}$ in $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$.
@@
We wish to apply Theorem 6.5 to demonstrate the claim. To this end, fix $\{\varphi_{i}\}_{i=1}^{k}\subset C_{c}(\mathbb{R}_{+}\times\mathcal{E}_{n})$ a finite collection of test functions. By Theorem 6.5 and the assumption that $\mathcal{L}^{N,n}_{G}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd}}}}\,\mathcal{L}$ we have that
@@
Indeed, by finiteness of the partition sizes it suffices to show this with the partition $\{(\{1\},\{2\})\}$. Observe that the number $U$ of selfing events in the ancestral line of the sample before the ancestral lines the sample experience an outcrossing satisfies
@@
At each selfing event, the two sample lineages have probability $\frac{1}{2}$ of coalescing. Therefore, the probability of not coalescing before the outcrossing is
@@
By Assumption 3.4 we know that $\alpha_{N}\to 1$, which shows that the non-coalescence probability of Equation (7.23) converges to 0 as $N\to\infty$. That $c_{N}T$ converges to 0 as $N\to\infty$ follows from Assumption 3.3 and the fact that the number of time-steps until a coalesce event or a dispersal is geometric with success probability $\frac{K_{N}}{N}\left(\frac{\alpha_{N}}{2}+1-\alpha_{N}\right)$. By a direct computation this is $O(c_{N})$ in expectation. The $\frac{K_{N}}{N}$ term comes from the probability that the individual containing the two sample lineages is a child, $\frac{\alpha_{N}}{2}$ is the probability that a selfing event occurs and yields a coalescence, and $1-\alpha_{N}$ is the probability of seeing an outcrossing.
We have established thus far that $\Pi_{G}^{N,n}(t)=\bar{\chi}^{N,n}(t)$ for all $t$ in which $S(\lfloor tc_{N}^{-1}\rfloor)\neq\lfloor tc_{N}^{-1}\rfloor$ with probability 1o(1)1-o(1). Further, we have shown that the measure of time $t$ on which $S(\lfloor tc_{N}^{-1}\rfloor)\neq\lfloor tc_{N}^{-1}\rfloor$ converges to 0 in distribution as $N$ goes to infinity. Therefore we have convergence in probability to 0 of Equation (7.22), as needed. ∎
@@
Notice by Lemma 7.13, it suffices to show convergence as random measures of $\mathcal{L}^{N,n}$ to $\mathcal{L}_{Q,\lambda}^{n}$. By Lemma 7.8 we have that $G^{N,n}$ converges weakly in $\mathcal{D}\left(\mathbb{R}_{+},\Theta\right)$ to a $Q$-$\lambda$ graph $G$. By Lemma 7.12 we know that the map $\Phi$ is continuous. Continuity preserves weak convergence and so $\mathcal{L}_{G}^{N,n}=\Phi(G^{N})$ converges weakly as a random measure to $\mathcal{L}_{Q,\lambda}^{n}=\Phi(G)$. By Lemma 7.14 we therefore have that $\mathcal{L}^{N,n}$ converges weakly as a random measure to $\mathcal{L}_{Q,\lambda}^{n}$. ∎
@@ Acknowledgements
Research supported by National Science Foundation grants DMS-2534011, DMS-2532674, DMS-2348164 and DMS-2152103. We appreciate Adam Jakubowski for pointing out references BK10a ; BK10b and the lack of method to construct a dense family of continuous bounded functions on the (non-locally compact) Skorokhod space.
Appendix Appendix A Convergence Criteria for the diploid model with selfing and overlapping generations
To provide convergence criteria for our diploid exchangeable model with selfing and overlapping generations, we proceed along the lines of [37] and [10, 1]. We define the offspring count $V_{i}$ for the $i$th individual to count the number of genetic descendents among the $2N$ genes in the next time-step that can trace their lineage back to the $i$th individual. Formally, we have
$V_{i}=\sum_{j}V_{i,j}+2V_{i,i}.$
@@
$V_{(1)}\geq V_{(2)}\geq\ldots\geq V_{(N)}$
the ranked version of the total offspring numbers $\left(V_{i}\right)_{1\leq i\leq N}$ and by
$\Phi_{N}:=\mathscr{L}\left(\frac{V_{(1)}}{2N},\frac{V_{(2)}}{2N},\ldots,\frac{V_{(N)}}{2N},0,0,\ldots\right)$
the law of the ranked offspring frequencies, viewed as a probability measure on the infinite ordered simplex $\Delta$.
Lemma Appendix A.1.
Suppose, as $N\to\infty$, that Assumption 3.3 holds and that
$\frac{1}{2c_{N}}\Phi_{N}(dx)\to\frac{1}{\left\langle x,x\right\rangle}\Xi^{\prime}(dx)$ (Appendix A.1)
vaguely on $\Delta\setminus\{\mathbf{0}\}$, where $\Xi^{\prime}$ is a probability measure on $\Delta$. Then Assumption 3.6 holds where, for each $n\geq 2$, $Q_{n}$ is the generator of an $n$-$\Xi$-coalescent with $\Xi=2\Xi^{\prime}$.
@@
Let $\mathit{P}_{N,n}=\left(\mathit{p}_{\xi\eta}^{N,n}\right)_{\xi,\eta\in\mathcal{E}_{n}}$ denote the one-step transition matrix of the ancestral process after applying the haploid map $F_{\rm hap}$ of Equation (3.10) so that
$\mathit{p}_{\xi\eta}^{N,n}=\mathbb{P}\left(F_{\rm hap}(\chi^{N,n}(1))=\eta\,\mid\,\chi^{N,n}(0)=\xi\right).$
To prove the claim we need to show that $c_{N}^{-1}\left(\mathit{P}_{N,n}-I\right)$ converges to a matrix $Q_{n}$ that is the generator of an $n$-$\Xi$-coalescent with $\Xi=2\Xi^{\prime}$.
@@
$Q_{N,n}=c_{N}^{-1}\left(\mathit{P}_{N,n}-I\right)=\Phi_{N}(A_{\varepsilon})\,Q_{N,n}^{\varepsilon}+\Phi_{N}(A_{\varepsilon}^{\mathrm{c}})\,Q_{N,n}^{\varepsilon,\mathrm{c}},$ (Appendix A.2)
where $Q_{N,n}^{\varepsilon}$ and $Q_{N,n}^{\varepsilon,\mathrm{c}}$ are the conditional transition matrices given $x\in A_{\varepsilon}$ and $x\in A_{\varepsilon}^{\mathrm{c}}$, respectively. By Assumption (Appendix A.1),
@@
and hence $\Phi_{N}(A_{\varepsilon})\to 0$ and $\Phi_{N}(A_{\varepsilon}^{\mathrm{c}})\to 1$ as $N\to\infty$ by Assumption 3.3.

### v2

@@ 7.3 Coalescing random walks on $\Omega$-valued processes
Note that $\Pi^{N,n}_{G}$ in Definition 7.9 is equal to $\Pi^{n}(\bar{G}^{N})$, where $\bar{G}^{N}$ is the sped-up ancestral graph of Lemma 7.8 and $\Pi^{n}$ is the map defined above.
@@
Let $\mathcal{E}_{\infty}:=\bigsqcup_{n\in\mathbb{N}}\mathcal{E}_{n}$ and recall the space $\Omega$ in (7.1). The measure-valued map $\Phi:\mathcal{D}\left(\mathbb{R}_{+},\Omega\right)\to\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{\infty}\right)\right)$ defined by
@@
Suppose that $g_{N}$ is a sequence of elements in $\mathcal{D}\left(\mathbb{R}_{+},\Omega\right)$ that converge therein to an element $g$. We need to show that $\Phi(g_{N})\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd\;}}}}\,\Phi(g)$. That is, we need to show that for any dense subset $\mathcal{F}$ of functions $f$ in $C_{b}(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{\infty}\right))$ that $\int fd\Phi(g_{N})$ converges to $\int fd\Phi(g)$. As $\Omega$ is a locally compact Polish space, by Theorem 6.5 it suffices to check that
@@
for any collection $\left(\varphi_{i}\right)_{1\leq i\leq k}$ of elements in $C_{c}\left(\mathbb{R}_{+}\times\Omega\right)$.
As $g_{N}$ converges to $g$ and $\Omega$ is a discrete space we have that, for large enough $N$, that $g_{N}(0)$ is eventually constant. In particular, we may restrict to the case that $g_{N}(0)=(n,m_{0},\xi_{0})$ for all $N$. This allows us to consider continuity of the restriction of $\Phi$ as a map into $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$ without loss of generality.
As each $\varphi_{i}$ is compactly supported on $\mathbb{R}_{+}\times\Omega$, there is a $T>0$ such that $\cup_{i}supp(\varphi_{i})\subset[0,T]\times\Omega$. Notice that, for any $N\geq 2$,
@@
The quantity on the right-hand-side of (7.20) depends only on discrete graph structure of $g_{N}$, i.e. the finite sequence of values that $g_{N}$ takes, and the edge lengths of the graph $g_{N}$, i.e. the jump times. As $\Omega$ is discrete and $g_{N}\to g$ in the Skorokhod topology, for large enough $N$ the discrete graph structure of $g_{N}$ is eventually constant. Further, the edge lengths of $g_{N}$ converge uniformly on $[0,T]$ to those of $g$. This yields the desired convergence.
@@
As a corollary, we have that if $G_{N}$ is a sequence of $\mathcal{D}\left(\mathbb{R}_{+},\Omega\right)$-valued random variables converging weakly to $G$, then $\Phi(G_{N})$ converges weakly in law to $\Phi(G)$. This is a key observation in the proof of Theorem 4.7.
@@
To prove Theorem 4.7, we establish a lemma demonstrating the equivalence of the quenched coalescent for the $Q$-$\lambda$ graph, as in Definition 7.3, and the quenched coalescent for an EFC, as in Definition 4.6. Before the formal statement of the lemma, we give an intuitive explanation for why the two constructions agree. Fix any finite sample of at most $n$ particles and follow their coalescing random-walk trajectories. Under both the $Q$-$\lambda$ graph and the EFC, coalescences of the walks occur with the same structures and rates, and fragmentations of each walk occur with the same rates. Because these rates uniquely determine the law of the induced coalescent on partitions and because both the $Q$-$\lambda$ graph and EFC give the same amount of information on the particle trajectories, the quenched coalescent laws coincide. The lemma that follows makes this argument rigorous via a joint coupling of the $Q$-$\lambda$ graph, the EFC, and random walks thereon.
@@
Suppose that $Q=\left(Q_{n}\right)_{n\in\mathbb{N}}$ is a consistent family of generators for a $\Xi$-coalescent for some measure $\Xi$ satisfying (3.11) (i.e. $2=\Xi(\Delta)$). Let $\Xi_{0}$ be the restriction of $\Xi$ on $\Delta\setminus\{\bf{0}\}$. Let $\Pi$ be an EFC process with characteristics $c_{k}=\Xi(\mathbf{0})$, $\nu_{Coag}=\nu_{\Xi_{0}}$, $c_{e}=\lambda$, and $\nu_{Disl}=0$. Then the random measures $\mathcal{L}_{\Pi}^{n}$ and $\mathcal{L}_{Q,\lambda}^{n}$ are equal in distribution.
@@
For a given $Q$-$\lambda$ graph $G=\left((l(t),m(t),\xi(t))\right)_{t\geq 0}$ and an EFC process $\Pi=\left(\Pi(t)\right)_{t\geq 0}$, we let $x^{G}=\left(x_{i}^{G}\right)_{1\leq i\leq n}$ and $x^{\Pi}=\left(x_{i}^{\Pi}\right)_{1\leq i\leq n}$ denote the family of coalescing random walks on $G$ and on $\Pi$, respectively. Let $\chi^{n}_{G}$ and $\chi^{n}_{\Pi}$ denote the associated coalescents for $x^{G}$ and $x^{\Pi}$, respectively. We describe a coupling of $\Pi$, $G$, $x^{\Pi}$, and $x^{G}$ together so that the claim holds. Specifically, it will be such that $\chi^{n}_{G}=\chi^{n}_{\Pi}$ pointwise, and such that $\mathbb{P}\left(\chi^{n}_{\Pi}\in\cdot\,\mid\,\Pi\right)=\mathbb{P}\left(\chi^{n}_{\Pi}\in\cdot\,\mid\,G\right)$, which together prove the claim. The construction follows that for the EFC (Ber, 04, Section 3.2).
@@
      and if $k(t)\in\{x_{i}^{\Pi}(t)\}_{1\leq i\leq n}$, then $G(t)=\left(l(t-),m(t),\xi(t-)\right)$, where $m(t)$ is equal to $x_{i}^{G}(t-)$ (for any $i$ such that $x_{i}^{\Pi}(t-)=k(t)$; check that this is indeed well-defined); for $H$ an independent Bernoulli random random variable with $\mathbb{P}\left(H=0\right)=\mathbb{P}\left(H=1\right)=\frac{1}{2}$ and if the only non-singleton block of $\xi^{(F)}(t)$ is $\{r,s\}$, then
@@
Suppose that, as $N\to\infty$, that $c_{N}\to 0$, $d_{N}c_{N}^{-1}\to\lambda<\infty$, and $\mathcal{L}^{N,n}_{G}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd\;}}}}\,\mathcal{L}$. Then $\mathcal{L}^{N,n}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd\;}}}}\,\mathcal{L}$.
@@
We apply Theorem 6.5 to demonstrate the claim. To this end, fix $\{\varphi_{i}\}_{i=1}^{k}\subset C_{c}(\mathbb{R}_{+}\times\mathcal{E}_{n})$ a finite collection of test functions. By Theorem 6.5 and the assumption that $\mathcal{L}^{N,n}_{G}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd\;}}}}\,\mathcal{L}$ we have that
@@
Indeed, by finiteness of the partition sizes it suffices to show this with the partition $\{(\{1\},\{2\})\}$. Observe that the number $U$ of selfing events in the ancestral line of the sample before the ancestral lines the sample experience an outcrossing event satisfies
@@
At each selfing event, the two sample lineages have probability $\frac{1}{2}$ of coalescing. Therefore, the probability of not coalescing before an outcrossing event is
@@
By Assumption 3.5 we know that $\alpha_{N}\to 1$, which shows that the non-coalescence probability of Equation (7.24) converges to 0 as $N\to\infty$. That $c_{N}T$ converges to 0 as $N\to\infty$ follows from Assumption 3.4 and the fact that the number of time-steps until a coalesce event or a dispersal is geometric with success probability $\frac{K_{N}}{N}\left(\frac{\alpha_{N}}{2}+1-\alpha_{N}\right)$. By a direct computation this is $O(c_{N})$ in expectation. The $\frac{K_{N}}{N}$ term comes from the probability that the individual containing the two sample lineages is an offspring, $\frac{\alpha_{N}}{2}$ is the probability that a selfing event occurs and yields a coalescence, and $1-\alpha_{N}$ is the probability of seeing an outcrossing event.
We have established thus far that $\Pi_{G}^{N,n}(t)=\bar{\chi}^{N,n}(t)$ for all $t\in(0,\infty)$ in which $S(\lfloor tc_{N}^{-1}\rfloor)\neq\lfloor tc_{N}^{-1}\rfloor$ with probability 1o(1)1-o(1). Further, we have shown that the measure of time $t$ on which $S(\lfloor tc_{N}^{-1}\rfloor)\neq\lfloor tc_{N}^{-1}\rfloor$ converges to 0 in distribution as $N\to\infty$. Therefore we have convergence in probability to 0 of Equation (7.23), as needed. ∎
@@
Notice by Lemma 7.13, it suffices to show weak convergence in law of $\mathcal{L}^{N,n}$ to $\mathcal{L}_{Q,\lambda}^{n}$. By Lemma 7.8 we have that $G^{N,n}$ converges weakly in $\mathcal{D}\left(\mathbb{R}_{+},\Omega\right)$ to a $Q$-$\lambda$ graph $G$. By Lemma 7.12 we know that the map $\Phi$ is continuous. Continuity preserves weak convergence and so $\mathcal{L}_{G}^{N,n}=\Phi(G^{N})$ converges weakly in law to $\mathcal{L}_{Q,\lambda}^{n}=\Phi(G)$. By Lemma 7.14 $\mathcal{L}^{N,n}$ converges weakly in law to $\mathcal{L}_{Q,\lambda}^{n}$. ∎
@@ Acknowledgements
Research supported by National Science Foundation grants DMS-2534011, DMS-2532574, DMS-2348164 and DMS-2152103. We appreciate Adam Jakubowski for pointing out references BK (10); Kou (16) and for discussion about convergence determining functions for measures on the Skorokhod space.
Appendix Appendix A Proof of the Theorem 3.11
We produce a lemma here to show that, for our notion of weak convergence in distribution that the intensity measure map sending a random measure $\mu$ to its intensity measure $\mathbb{E}\left[\mu\right]$ is continuous.
Lemma Appendix A.1.
Let $\left(\mu_{N}\right)_{N\in\mathbb{N}}$ be a sequence of random variables taking value in $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$, where $E$ is a locally compact Polish space. Suppose that $\mu_{N}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd\;}}}}\,\mu$ in $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$. Then the sequence of intensity measures $\mathbb{E}\left[\mu_{N}\right]\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\mathbb{E}\left[\mu\right]$ in $\mathbb{R}$.
Proof.
Recall that for a random measure $\mu$ the intensity measure $\mathbb{E}\left[\mu\right]$ is defined as the unique measure for which
$\mathbb{E}\left[\mu\right](f):=\mathbb{E}\left[\mu(f)\right]$
for any $f$ in $C_{b}(\mathcal{D}\left(\mathbb{R}_{+},E\right))$.
We need to show that
$\mathbb{E}\left[\mu_{N}(f)\right]=\mathbb{E}\left[\mu_{N}\right](f)\to\mathbb{E}\left[\mu\right](f)=\mathbb{E}\left[\mu(f)\right]$
for any $f$ in $C_{b}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$. Fix any such $f$. The topology on $\mathcal{M}_{1}(\mathcal{D}\left(\mathbb{R}_{+},E\right))$ is such that the evaluation maps $T_{f}$ sending $\mu$ to $\mu(f)$ is continuous and bounded. As $\mu_{N}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd\;}}}}\,\mu$ we have that
$T_{f}(\mu_{N})=\mu_{N}(f)\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,T_{f}(\mu)=\mu(f).$
As $T_{f}$ is continuous and bounded, convergence in distribution implies convergence in expectation. This yields the claim.
With Theorem 4.7 in hand, we are ready to prove Theorem 3.11 as a corollary.
Proof of Theorem 3.11.
Note that the assumptions of Theorem 3.11 are the same as those of Theorem 4.7. Consequently, by Theorem 4.7 we have that the sequence of random measures $\mathcal{L}^{N,n}$ converges weakly in law to $\mathcal{L}_{\Pi}^{n}$ where $\Pi$ is an EFC with coalescence measure $\Xi$ on $\Delta$, no non-binary fragmentation, and binary fragmentation with rate $\lambda$. By Lemma Appendix A.1, the intensity measure mapping is continuous, and so $\mathbb{E}\left[\mathcal{L}^{N,n}\right]=\mathbb{P}_{\xi_{0}^{n}}\left(\bar{\chi}^{N,n}\in\cdot\right)$ converges in distribution to $\mathbb{E}\left[\mathcal{L}_{\Pi}^{n}\right]$. It suffices, therefore, to show that $\mathbb{E}\left[\mathcal{L}_{\Pi}^{n}\right]$ is the law of an $n$-$\Xi$-coalescent.
Appendix Appendix B Convergence criteria for the general model
To provide convergence criteria for our diploid exchangeable model with selfing and overlapping generations, we proceed along the lines of [35] and [9, 1]. We define the genetic contribution $\tilde{V}_{i}$ for the $i$th individual to count the number of genetic descendents among the $2N$ genes in the next time-step that can trace their lineage back to the $i$th individual. Formally, we have
$\tilde{V}_{i}=\sum_{j}V_{i,j}+V_{i,i}.$
@@
$\tilde{V}_{(1)}\geq\tilde{V}_{(2)}\geq\ldots\geq\tilde{V}_{(N)}$
the ranked version of the total offspring numbers $\left(\tilde{V}_{i}\right)_{1\leq i\leq N}$ and by
$\Phi_{N}:=\mathfrak{L}\left(\mathfrak{V}_{N}\right)$
the law on $\Delta$ of the ranked offspring frequencies $\mathfrak{V}_{N}$ defined by
$\mathfrak{V}_{N}:=\left(\frac{\tilde{V}_{(1)}}{2N},\frac{\tilde{V}_{(2)}}{2N},\ldots,\frac{\tilde{V}_{(N)}}{2N},0,0,\ldots\right).$ (Appendix B.1)
Remark Appendix B.1.
We recall here the transition rate $q^{n}(\xi,\eta)=\lambda_{b;k_{1},\ldots,k_{r};s}$ from a state $\xi$ to a state $\eta$ for an $n$-$\Xi$-coalescent, where $\xi$ consists of $b$ blocks and $\eta$ is of $(k_{1},\ldots,k_{r};s)$-type with respect to $\xi$. That is, $\eta$ is obtained from $\xi_{0}$ by keeping $s$ of the blocks the same and coalescing the remaining bsb-s blocks into $r$ blocks made up of $k_{1},\ldots,k_{r}$ blocks from $\xi$. By [9, Equation 1.9]
$q^{n}(\xi,\eta)=\lambda_{b;k_{1},\ldots,k_{r};s}$ $=\mathbbm{1}_{r=1;k_{1}=2}\Xi(\mathbf{0})$ (Appendix B.2)
$+\int_{\Delta\setminus\{\mathbf{0}\}}\sum_{l=0}^{s}\sum_{i_{1},\ldots,i_{r}\ \mathrm{distinct}}\binom{s}{l}\Big(\prod_{j=1}^{r}x_{i_{j}}^{k_{j}}\Big)\left(\prod_{j=r+1}^{l}x_{i_{j}}\right)\,(1-|x|)^{s-l}\frac{\Xi(dx)}{\left\langle x,x\right\rangle}.$ (Appendix B.3)
Lemma Appendix B.2.
Suppose that Assumption 3.4 holds and that, as $N\to\infty$,
$\frac{1}{2c_{N}}\Phi_{N}(dx)\to\frac{1}{\left\langle x,x\right\rangle}\Xi^{\prime}(dx)$ (Appendix B.4)
vaguely on $\Delta\setminus\{\mathbf{0}\}$, where $\Xi^{\prime}$ is a probability measure on $\Delta$ (i.e. $\Xi^{\prime}(\mathbf{0})=1-\Xi^{\prime}(\Delta)$). Then Assumption 3.7 holds where, for each $n\geq 2$, $Q_{n}$ is the generator of an $n$-$\Xi$-coalescent with $\Xi=2\Xi^{\prime}$.
@@
Let $\mathit{H}_{N,n}=\left(h_{\xi\eta}^{N,n}\right)_{\xi,\eta\in\mathcal{E}_{n}}$ denote the one-step transition matrix of the ancestral process after applying the haploid map $F_{\rm hap}$ of Equation (3.9) so that
$h_{\xi\eta}^{N,n}=\mathbb{P}\left(F_{\rm hap}(\chi^{N,n}(1))=\eta\,\mid\,\chi^{N,n}(0)=\xi\right).$
To prove the claim we need to show that $\frac{1}{c_{N}}\left(\mathit{H}_{N,n}-I\right)$ converges to a matrix $Q_{n}$ that is the generator of an $n$-$\Xi$-coalescent with $\Xi=2\Xi^{\prime}$.
@@
$Q_{N,n}:=\frac{1}{c_{N}}\left(\mathit{H}_{N,n}-I\right)=\Phi_{N}(A_{\varepsilon})\,Q_{N,n}^{\varepsilon}+\Phi_{N}(A_{\varepsilon}^{\mathrm{c}})\,Q_{N,n}^{\varepsilon,\mathrm{c}},$ (Appendix B.5)
where $Q_{N,n}^{\varepsilon}$ and $Q_{N,n}^{\varepsilon,\mathrm{c}}$ are the conditional transition rate matrices given the events $\{\mathfrak{V}_{N}\in A_{\varepsilon}\}$ and $\{\mathfrak{V}_{N}\in A_{\varepsilon}^{\mathrm{c}}\}$ respectively, where $\mathfrak{V}_{N}$ is the ranked offspring frequencies (Appendix B.1). That is,
$Q_{N,n}^{\varepsilon}$ $=\frac{1}{c_{N}}\left(\mathbb{P}\left(F_{\rm hap}(\chi^{N,n}(1))=\eta\,\mid\,\chi^{N,n}(0)=\xi,\mathfrak{V}_{N}\in A_{\varepsilon}\right)-\delta_{\xi\eta}\right)\text{ and }$
$Q_{N,n}^{\varepsilon,c}$ $=\frac{1}{c_{N}}\left(\mathbb{P}\left(F_{\rm hap}(\chi^{N,n}(1))=\eta\,\mid\,\chi^{N,n}(0)=\xi,\mathfrak{V}_{N}\in A_{\varepsilon}^{c}\right)-\delta_{\xi\eta}\right).$
By Assumption (Appendix B.4),
@@
and hence $\Phi_{N}(A_{\varepsilon})\in O_{\varepsilon}(c_{N})$ and $\Phi_{N}(A_{\varepsilon}^{\mathrm{c}})\to 1$ as $N\to\infty$ by Assumption 3.4. Here $O_{\varepsilon}(1)$ simply means that, the term is $O(1)$ for any fixed $\varepsilon>0$.